### v1

@@ 5 On the semismooth property of solution maps to inclusions
Due to the definition of the SCD property around a point and Proposition 5.10 we can find an open neighborhood $U^{\prime}\times V^{\prime}\times W^{\prime}$ of $(\bar{x},\bar{y},0)$ such that for every $(x,y,z)\in U^{\prime}\times V^{\prime}\times W^{\prime}$ the mapping $F$ has the SCD property around $(x,y,z)$ and ${\rm scd\,reg\,}F(x,y,z)<\infty$. Next consider the open neighborhood $U^{\prime\prime}:=\{x\in\Omega\cap U^{\prime}\,\big{|}\,\sigma(x)\in V^{\prime}\}$ and choose an open neighborhood $U$ of $\bar{x}$ satisfying $\mathrm{cl}\,U\subset U^{\prime\prime}$. This choice of $U$ ensures that ${\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$, $x\in U$. Next we show that $\Psi$ is locally bounded at $x\in U$. Indeed, we have $\gamma:={\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$ and, due to the continuity of $\sigma$ and Proposition 5.10, we can find a neighborhood $\tilde{U}\subset\Omega$ of $x$ satisfying ${\rm scd\,reg\,}F(x^{\prime},\sigma(x^{\prime}),0)<\gamma+1$ $\forall x^{\prime}\in\tilde{U}$. Hence, $\|X^{*}\|<\gamma+1$ for every $X^{*}\in\Psi(x^{\prime})$ and our claim is established. In order to show the last assertion, consider $x\in U$ such that $F$ is SCD semismooth at $(x,\sigma(x),0)$, set $\gamma:={\rm scd\,reg\,}F(x,\sigma(x),0)$ and choose $\tilde{\delta}>0$ such that ${\rm scd\,reg\,}F(x^{\prime},\sigma(x^{\prime}),0)<\gamma+1$ $\forall x^{\prime}\in{\cal B}_{\tilde{\delta}}(x)$. By definition of the SCD semismooth property we can find for every $\epsilon>0$ some $\delta\in(0,\tilde{\delta})$ such that for every $x^{\prime}\in{\cal B}_{\delta}(x)$, every subspace $L^{*}\in{\mathcal{S}}^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ and every $p\in\mathbb{R}^{m}$ we have
@@
and $\kappa(L^{*})\leq\gamma+1$, showing
@@
with $\gamma^{\prime}=\sqrt{1+(1+\gamma)^{2}}$. Thus $(1-\epsilon\gamma^{\prime})\|\sigma(x^{\prime})-\sigma(x)\|\leq\|x^{\prime}-x\|(\|X_{L^{*}}^{T}\|+\epsilon\gamma^{\prime})$ implying $\|\sigma(x^{\prime})-\sigma(x)\|\leq 2\|x^{\prime}-x\|((\gamma+1)+\frac{1}{2})$ whenever $\epsilon\in(0,1/(2\gamma^{\prime}))$. Hence
@@
Consider a mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ and a continuous mapping $\sigma:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, satisfying $0\in F(x,\sigma(x))$, $x\in\Omega$. Further assume that for every $x\in\Omega$ the mapping $F$ is SCD semismooth at $(x,\sigma(x),0)$ and $\mathrm{scd\,reg}\,F(x,\sigma(x),0)<\infty$. Then $\sigma$ is semismooth on $\Omega$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n\times m}$ given by
$\Psi(x):=\begin{cases}\{-X_{L^{*}}^{T}\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\}&\mbox{if $x\in\Omega$},\\ \emptyset&\mbox{else.}\end{cases}$
@@
$\Theta(x):=\begin{cases}\{g_{x}^{T}-g_{y}^{T}X_{L^{*}}^{T}\,\big{|}\,(g_{x}^{T},g_{y}^{T})\in\Phi(x,\sigma(x)),L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\}&\mbox{if $x\in\Omega$},\\ \emptyset&\mbox{else.}\end{cases}$
Note that in Corollary 5.5 we require that $\sigma$ is some continuous localization of the solution map $S$, whereas in Corollary 5.12 the mapping $\sigma$ only needs to be any continuous selection of $S$. This is illustrated by the following academic example.
@@
Consider the mapping $F:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ given by $F(x,y)=|y|-|x|$. Then the solution map $S$ to the equation $F(x,y)=0$ is given by
$S(x)=\{x\}\cup\{-x\}$
and no single-valued localization of $S$ around the origin exists. On the other hand, consider a continuous selection $\sigma$, e.g., $\sigma(x)=\{x\}$. It is easy to compute by means of Lemma 2.13 that
${\mathcal{S}}^{*}F(x,\sigma(x),0)=\begin{cases}\{{\rm rge\;}\big{(}1,(-1,1)\big{)}\}&\mbox{if $x\not=0$,}\\ \{{\rm rge\;}\big{(}1,(-1,1)\big{)},{\rm rge\;}\big{(}1,(1,1)\big{)},{\rm rge\;}\big{(}1,(-1,-1)\big{)},{\rm rge\;}\big{(}1,(1,-1)\big{)}\}&\mbox{if $x=0$.}\end{cases}$
We see that all subspaces $L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)$, $x\in\mathbb{R}$ belong to ${\cal Z}_{1,2}^{\rm reg}$ and therefore $\sigma$ is semismooth on $\mathbb{R}$ with respect to the mapping
$\Psi(x)=\begin{cases}\{1\}&\mbox{if $x\not=0$,}\\ \{1,-1\}&\mbox{if $x=0$.}\end{cases}$
We see that $\nabla\sigma(x)\subset\Psi(x)$, $x\in\mathbb{R}$, but at the origin $\Psi(0)$ also contains the value $-1$ which has nothing to do with the derivative of $\sigma$. Note that we can arbitrarily modify $\Psi$ on a finite subset of $\mathbb{R}$ without destroying the semismooth property.
@@ 6 Conclusion
The paper contains a deep analysis of some generalizations of the classical semismoothness property. Namely, it deals with the so-called $\Psi$-semismoothness and new notions of $\Gamma$-semismoothness and SCD-semismoothness. These notions are closely related to each other and have a number of useful properties concerning, in particular, some fundamental numerical approaches to nonsmooth problems. For instance, whereas the standard rules of generalized differential calculus attain typically the form of inclusions, the presence of the mentioned semismoothness properties enables us to construct suitable approximations of the considered mappings which fulfill the respective calculus rules as equalities. This is especially helpful in connection with a family of bundle methods in nonsmooth optimization.
A special attention has been paid to the semismoothness of localizations (selections) of solution maps to parameter-dependent inclusions and their possible compositions with real-valued objectives. The presented analysis of these mappings paves the way to an efficient usage of the ImP approach in a broad class of problems with equilibrium constraints.

### v2

@@ 5 On the semismooth property of solution maps to inclusions
Due to the definition of the SCD property around a point and Proposition 5.8 we can find an open neighborhood $U^{\prime}\times V^{\prime}\times W^{\prime}$ of $(\bar{x},\bar{y},0)$ such that for every $(x,y,z)\in U^{\prime}\times V^{\prime}\times W^{\prime}$ the mapping $F$ has the SCD property around $(x,y,z)$ and ${\rm scd\,reg\,}F(x,y,z)<\infty$. Next consider the open neighborhood $U^{\prime\prime}:=\{x\in\Omega\cap U^{\prime}\,\big{|}\,\sigma(x)\in V^{\prime}\}$ and choose an open neighborhood $U$ of $\bar{x}$ satisfying $\mathrm{cl}\,U\subset U^{\prime\prime}$. This choice of $U$ ensures that ${\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$, $x\in U$. We now show that $\Psi$ is locally bounded at $x\in U$. Indeed, we have $\gamma:={\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$ and, due to the continuity of $\sigma$ and Proposition 5.8, we can find a neighborhood $\tilde{U}\subset\Omega$ of $x$ satisfying ${\rm scd\,reg\,}F(x^{\prime},\sigma(x^{\prime}),0)<\gamma+1$ $\forall x^{\prime}\in\tilde{U}$. Hence, $\|X^{*}\|<\gamma+1$ for every $X^{*}\in\Psi(x^{\prime})$ and our claim is established. In order to show the last assertion, consider $x\in U$ such that $F$ is SCD semismooth at $(x,\sigma(x),0)$, set $\gamma:={\rm scd\,reg\,}F(x,\sigma(x),0)$ and choose $\tilde{\delta}>0$ such that ${\rm scd\,reg\,}F(x^{\prime},\sigma(x^{\prime}),0)<\gamma+1$ $\forall x^{\prime}\in{\cal B}_{\tilde{\delta}}(x)$. By definition of the SCD semismooth property we can find for every $\epsilon>0$ some $\delta\in(0,\tilde{\delta})$ such that for every $x^{\prime}\in{\cal B}_{\delta}(x)$, every subspace $L^{*}\in{\mathcal{S}}^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ and every $p\in\mathbb{R}^{m}$ we have
@@
and $\kappa(L^{*})\leq\gamma+1$. This shows that
@@
with $\gamma^{\prime}=\sqrt{1+(1+\gamma)^{2}}$. Thus $(1-\epsilon\gamma^{\prime})\|\sigma(x^{\prime})-\sigma(x)\|\leq\|x^{\prime}-x\|(\|X_{L^{*}}^{T}\|+\epsilon\gamma^{\prime})$ implying $\|\sigma(x^{\prime})-\sigma(x)\|\leq 2\|x^{\prime}-x\|((\gamma+1)+1/2)$ whenever $\epsilon\in(0,1/(2\gamma^{\prime}))$. Hence
@@
Consider a mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ and a continuous mapping $\sigma:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, satisfying $0\in F(x,\sigma(x))$, $x\in\Omega$. Further assume that for every $x\in\Omega$ the mapping $F$ is SCD semismooth at $(x,\sigma(x),0)$ and $\mathrm{scd\,reg}\,F(x,\sigma(x),0)<\infty$. Then $\sigma$ is semismooth on $\Omega$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ given by
$\Psi(x):=\{-X_{L^{*}}^{T}\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\},\ x\in\mathrm{dom}\,\Psi:=\Omega.$
@@
$\Theta(x):=\{g_{x}^{T}-g_{y}^{T}X_{L^{*}}^{T}\,\big{|}\,(g_{x}^{T},g_{y}^{T})\in\Phi(x,\sigma(x)),L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\},\ x\in\mathrm{dom}\,\Theta:=\Omega.$
Note that both in Corollary 5.5 and Corollary 5.10 the $\Psi$-semismoothness of $\sigma$ on $\Omega$ implies that $\sigma$ is locally Lipschitz and almost everywhere differentiable on $\Omega$. However, in Corollary 5.5 we require that $\sigma$ is some continuous localization of the solution map $S$, whereas in Corollary 5.10 the mapping $\sigma$ only needs to be any continuous selection of $S$. In case of a single-valued mapping $F$, a similar result has already been stated in [21]. However, since most implicit function theorems (see, e.g., [16, 25, 31]) guarantee existence of a locally unique implicit function, it is clear that the assumptions of such theorems cannot be satisfied in situations where $\sigma$ is not locally unique.
To illustrate Corollary 5.10 consider the following academic example.
@@
Consider the functions $f_{1},f_{2}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and the mapping $F:\mathbb{R}\times\mathbb{R}\rightrightarrows\mathbb{R}$ given by
$f_{1}(x,y):=\frac{1}{2}y^{2}-xy,\ f_{2}:=-\frac{1}{2}y^{2},\ F(x,y):=\partial^{c}_{y}f(x,y)\ \mbox{with}\ f(x,y):=\max\{f_{1}(x,y),f_{2}(x,y)\}.$ (5.21)
[…]
$f(x,y)=\begin{cases}f_{2}(x,y)&\mbox{for $y\in[x^{-},x^{+}]$},\\ f_{1}(x,y)&\mbox{for $y\not\in(x^{-},x^{+})$,}\end{cases}$
where $x^{-}:=\min\{x,0\}$ and $x^{+}:=\max\{x,0\}$. Thus, for any $x\in\mathbb{R}$ we obtain
$\mathrm{gph}\,\partial^{c}_{y}f(x,\cdot)$ =$=\{(y,y-x)\,\big{|}\,y\not\in[x^{-},x^{+}]\}\cup\{(y,-y)\,\big{|}\,y\in(x^{-},x^{+})\}$
($\qquad\cup(\{x^{-}\}\times[x^{-}-x,-x^{-}])\cup(\{x^{+}\}\times(-x^{+},x^{+}-x))$
and we see that $S(x)=\{x,0\}$ does not possess a single-valued localization around $(0,0)$. Let us consider the continuous selection $\sigma(x)=\{x\}=\mathop{\rm arg\,min}_{y}f(x,y)$. We compute now the SC derivative ${\mathcal{S}}F(x,x,0)$ for $x\in\mathbb{R}$. By the above representation of $\mathrm{gph}\,\partial^{c}_{y}f(x,\cdot)$ it follows that ${\cal O}_{F}$ is the union of the four sets
$O_{1}:=\{(x,y,y-x)\,\big{|}\,x\in\mathbb{R},y\not\in[x^{-},x^{+}]\},\ O_{2}:=\{(x,y,-y)\,\big{|}\,x\not=0,y\in(x^{-},x^{+})\}$
$O_{3}:=\{(x,x,z^{*})\,\big{|}\,x\not=0,z^{*}\in(x^{-}-x,x^{+}-x)\},\ O_{4}:=\{(x,0,z^{*})\,\big{|}\,x\not=0,z^{*}\in(-x^{+},-x^{-})\}.$
Moreover, defining the four subspaces
$L_{1}:=\{(u,v,v-u)\,\big{|}\,(u,v)\in\mathbb{R}^{2}\},\ L_{2}:=\{(u,v,-v)\,\big{|}\,(u,v)\in\mathbb{R}^{2}\},$
$L_{3}:=\{(u,u,v)\,\big{|}\,(u,v)\in\mathbb{R}^{2}\},\ L_{4}:=\{(u,0,v)\,\big{|}\,(u,v)\in\mathbb{R}^{2}\},$
we have that $T_{\mathrm{gph}\,F}(x,y,z^{*})=L_{k}$ for all $(x,y,z^{*})\in O_{k}$ and $k=1,\ldots,4$. Hence, ${\mathcal{S}}F(x,\sigma(x),0)=\{L_{1},L_{3}\}$ for $x\not=0$ and ${\mathcal{S}}F(0,0,0)=\{L_{1},L_{2},L_{3},L_{4}\}$. All the adjoint subspaces
$L_{1}^{*}={\rm rge\;}(1,-1,1),\ L_{2}^{*}={\rm rge\;}(-1,0,1),\ L_{3}^{*}={\rm rge\;}(0,-1,1),\ L_{4}^{*}={\rm rge\;}(0,0,1)$
belong to ${\cal Z}_{1,2}^{\rm reg}$ and thus we can conclude from Corollary 5.10 that $\sigma$ is semismooth with respect to
$\Psi(x):=\begin{cases}\{1\}&\mbox{if $x\not=0$,}\\ \{1,0\}&\mbox{if $x=0$.}\end{cases}$ (5.22)
Note that in accordance with Theorem 3.8(iii) the mapping $\Psi$ is almost everywhere single-valued, although ${\mathcal{S}}F(x,\sigma(x),0)$ does not have this property for any $x$.
These results enable us to apply the considered ImP approach to the following nonsmooth bilevel program
$\min_{(x,y)\in\mathbb{R}^{2}\times\mathbb{R}}\varphi(x,y):=2\big{|}y-|x_{1}|\big{|}-|x_{2}|+\frac{1}{2}x_{1}^{2}\quad\mbox{subject to}\quad y\in\mathop{\rm arg\,min}f(\eta(x),\cdot),$
where $f$ is given by (5.21) and the control variable $x$ enters now the lower-level problem via the nonsmooth function $\eta:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ given by $\eta(x)=|x_{1}|-|x_{2}|$. The solution of this bilevel program is $\bar{x}=(0,0)$, $\bar{y}=0$ and we solved it by applying the BT algorithm to the (reduced) problem
$\min\theta(x):=\varphi(x,\sigma(\eta(x))),$
where $\sigma(\xi):=\mathop{\rm arg\,min}_{y}f(\xi,y)$ is evaluated by solving the nonlinear program
$\min_{z,y}z\quad\mbox{subject to}\quad z\geq f_{1}(\xi,y),\ z\geq f_{2}(\xi,y)$
by means of the MATLAB routine fmincon, which was able to compute in every case the correct global solution. Since ||$\cdot$|| ⋅ | is G-semismooth, i.e., |$\partial$|$\cdot$|∂ | ⋅ |-ss, and $\sigma$ is $\Psi$-ss with $\Psi$ given by (5.22), these functions are also semismooth with respect to the selections
$t\mapsto{\rm sgn}(t):=\begin{cases}1&\mbox{if $t\geq 0$,}\\ -1&\mbox{for $t<0$,}\end{cases}\quad\xi\mapsto\tilde{\Psi}(\xi):=\begin{cases}1&\mbox{if $\xi\not=0$,}\\ 0&\mbox{if $\xi=0$}\end{cases}$
of |$\partial$|$\cdot$|∂ | ⋅ | and $\Psi$, respectiveley. By means of Lemma 3.2 and Corollary 5.10 we obtain that $\theta$ is semismooth with respect to the mapping
$(x_{1},x_{2})\mapsto\big{(}2{\rm sgn}(u(x))(\tilde{\Psi}(\eta(x))-1){\rm sgn}(x_{1})+x_{1},-2{\rm sgn}(u(x))\tilde{\Psi}(\eta(x)){\rm sgn}(x_{2})-{\rm sgn}(x_{2})\big{)},$
where $u(x):=\sigma(\eta(x))-|x_{1}|$. We did not make any attempt to compute exact Clarke subgradients of $\theta$ because, according to our theory, elements of a suitable semismooth derivative may replace them very well. Indeed, starting from ${x}^{(0)}=(5,-1)$, the BT algorithm stopped after $16$ Iterations and $21$ evaluations of the oracle at the point $(-2.8\cdot 10^{-6},1.8\cdot 10^{-8})$. Other starting points yield similar results.
@@ 6 Conclusion
The paper contains a deep analysis of some generalizations of the classical semismoothness property. Namely, it deals with the so-called $\Psi$-semismoothness and new notions of $\Gamma$-semismoothness and SCD-semismoothness. These notions are closely related to each other and have a number of useful properties concerning, in particular, some fundamental numerical approaches to nonsmooth problems. For instance, whereas the standard rules of generalized differential calculus attain typically the form of inclusions, the calculus of semismooth derivatives yields again a semismooth derivative. This is especially helpful in connection with a family of bundle methods in nonsmooth optimization.
A special attention has been paid to the semismoothness of localizations and selections of solution maps to parameter-dependent inclusions and their possible compositions with real-valued objectives. The presented analysis of these mappings paves the way to an efficient extension of the ImP approach to a broader class of problems with equilibrium constraints.
Acknowledgment
The authors would like to thank both reviewers for their numerous helpful comments improving the quality of the paper.